. A 0.01 M HCl solution is diluted with water hundred times. 1. The pH of the solution increases by 2. 2. The pOH of the solution increases by 2. 3. The hydronium ion concentration of the solution decreases from10─2 M to10─4 M. 4. The hydroxide ion concentration of the solution does not change.


1
Expert's answer
2022-05-16T17:02:10-0400

Solution:

HCl is strong electrolyte. It completely dissociates.

HCl(aq) + H2O = H3O+(aq) + Cl(aq)

Thus, its hydronium ion concentration [H3O+] will be 0.01 M

[H3O+] = 0.01 M (before dilution)

The pH of the solution will be −log[H3O+] = −log(0.01) = 2

pH = 2 (before dilution)

pOH = 14 − pH = 14 − 2 = 12

pOH = 12 (before dilution)


When 0.01 M HCl solution is diluted 100 times, the molarity of solution will be (0.01 / 100) = 0.0001 M

Thus, its hydronium ion concentration [H3O+] will be 0.0001 M

[H3O+] = 0.0001 M (after dilution)

The pH of the solution will be −log[H3O+] = −log(0.0001) = 4

pH = 4 (after dilution)

pOH = 14 − pH = 14 − 4 = 10

pOH = 10 (after dilution)


Correct options are 1 and 3

1) The pH of the solution increases by 2

3) The hydronium ion concentration of the solution decreases from10−2 M to10−4 M